OpenStudy (anonymous):

2. Henry can clean the pool three times as fast as John. Together they can do the job in 2 hours. About how long would it take Henry to clean the pool alone?

OpenStudy (anonymous):

Henry = t/3 John = t Together they can do the job in 2 hours t/3 + t = 2 hours 4t/3 = 2 hours t =3/2 hours Henry = t/3 = (3/2)/3 = 0.5 hours

OpenStudy (anonymous):

thank you!